OpenStudy (anonymous):

What is the equation of a line that passes through (-6, 2) and has a slope of −12 ? A) y=x2−2 B) y=−x2−1 C) y=2x+1 D) y=−2x−1

OpenStudy (muzzack):

could you fan me ;)

OpenStudy (anonymous):

I did lol

OpenStudy (muzzack):

y-2=-12(x+6) y-2=-12x-72 y=-12x-70

OpenStudy (anonymous):

crap. in the answers its x/2 for the first two. and the last two its the same